OpenStudy (1645323):

The polynomial 4x3 -8x2 -11x - 3 represents the volume in cubic feet of a rectangular fish tank,. The area of the base can be expressed by the polynomial 4x2 + 4x + 1. Find an expression for the height of the fish tank. Select one: a. 2x + 1 b. x - 3 c. x + 3 d. x2 + 2x + 1

OpenStudy (1645323):

@terenzreignz

OpenStudy (1645323):

@mathmale

OpenStudy (mathmale):

Hello, 1645323! Recall that the volume of a rectangular 3D figure is (length) * (width) * ( height). Supposing the volume is V = L*W*H, how would you solve this equation for H?

OpenStudy (mathmale):

Hint: divide both sides of this equation by what quantity?

OpenStudy (1645323):

the area of the base?

OpenStudy (mathmale):

Right. In this problem you are given a polynomial expression which represents the volume of a rectangular prism, along with one that represents the area of the base. Using my previous question as a guide, how would you determine the height of this rectangular prism?

OpenStudy (mathmale):

V = 4x3 -8x2 -11x - 3 Area of base = 4x2 + 4x + 1 A formula for the height is ... ?

OpenStudy (1645323):

idk what is it?@mathmale

OpenStudy (mathmale):

Take V and divide it by the Area of the Base: 4x3 -8x2 -11x - 3 H = ------------------ 4x2 + 4x + 1 You could use long division here. An alternative approach would be synthetic division. Use whichever is the more familiar to you.

OpenStudy (1645323):

its x-3

OpenStudy (mathmale):

____________________ 4x^2+4x+1 | 4x3 -8x2 -11x - 3 Yes, that's x-3. Congrats! Nice work!